# Improving Autoformalization Using Direct Dependency Retrieval

### v2

@@ Abstract
Shaoqi Wang† Lu Yu‡ Siwei Lou† Feng Yan∗ Chunjie Yang† † Zhejiang University, {sq_w,swlou,cjyang999}@zju.edu.cn ‡ Ant Group ∗ Hong Kong Polytechnic University, {feng555.yan}@polyu.edu.hk
@@ 4 Direct Dependency Retrieval-augmented Autoformalization
Method Diff01 Diff23 Diff45 Diff67 Diff89 Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Claude-3.5-Sonnet DeepSeek-R1 GPT-4o Ling-flash-2.0 Qwen-Max R@5 R@10 DDR 0.85 0.74 0.92 0.58 0.91 0.68 0.91 0.57 0.82 0.44
@@ 4.4 Retrieval Results
After filtering, the results show substantial performance disparities among the models. Despite having no hallucinations, R@5 and R@10 yield the lowest precision (0.02-0.10) and recall (0.10-0.20), indicating their retrieval capability is severely limited. The filtered prompt-based methods achieve moderate performance, with precision scores ranging from 0.2 to 0.5 and recall scores from 0.3 to 0.5. Our proposed DDR method demonstrates the most superior performance, achieving a high precision of 0.8-0.9 and a recall of 0.4-0.8, significantly outperforming all other baseline methods.

### v3

@@ Abstract
Shaoqi Wang† Lu Yu‡ Siwei Lou† Feng Yan∗ Chunjie Yang† Qing Cui‡ Jun Zhou‡ † Zhejiang University, ‡ Ant Group, ∗ Hong Kong Polytechnic University
@@ 4 Direct Dependency Retrieval-augmented Autoformalization
Method Diff01 Diff23 Diff45 Diff67 Diff89 Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Claude-3.5-Sonnet DeepSeek-R1 GPT-4o Ling-flash-2.0 Qwen-Max R@5 R@10 DDR 0.84 0.82 0.90 0.87 0.91 0.92 0.91 0.88 0.83 0.83
@@ 4.4 Retrieval Results
After filtering, the results show substantial performance disparities among the models. Despite having no hallucinations, R@5 and R@10 yield the lowest precision (0.02-0.10) and recall (0.10-0.20), indicating their retrieval capability is severely limited. The filtered prompt-based methods achieve moderate performance, with precision scores ranging from 0.2 to 0.5 and recall scores from 0.3 to 0.5. Our proposed DDR method demonstrates the most superior performance, achieving a high precision of 0.8-0.9 and a recall of 0.8-0.9, significantly outperforming all other baseline methods.
@@ C.2 Metrics
C.3 Hallucination Details
It is important to note that the hallucination metric, Hall, is computed exclusively over the set of effective samples. We define an effective sample as an instance for which the model retrieves one or more dependencies from the library. Some autoformalization tasks are simple enough to be solved using only the given context and basic syntax, thereby not requiring any explicit library dependencies. For such cases, our model is designed to output a predefined token, "None" (or "No usage" in our implementation), to indicate that no dependency retrieval is necessary. Since these instances do not involve an actual retrieval action, the concept of hallucination, retrieving an incorrect or non-existent item, is not applicable. Therefore, they are excluded from the metric’s calculation. For example, in a dataset of 200 retrieved dependency samples, if 50 are resolved with a "None" output, the effective sample size for computing the mean hallucination rate (Hall) and its standard deviation (Hall) is 150.
C.4 Baseline Details